OpenStudy (anonymous):

A block of mass 5 kg is moving horizontally at a speed of 1.5 m/s. A perpendicular force of 5N acts on it for 4 sec. What will be the distance of the block from the point where the force started acting. A. 10m. B. 8m C.6m. D.2m

OpenStudy (anonymous):

@irishboy123

OpenStudy (anonymous):

@arindameducationusc

OpenStudy (anonymous):

@michele_laino

OpenStudy (michele_laino):

the situation of your problem can be summarized as follows:  namely the motion of the block is a motion which is contained inside the horizontal plane \(x,y\). Gravity is acting perpendicularly with respect to that plane, therefore, it is represented by a vector which is not contained inside such motion plane. Now on the block is acting the external force \(F\), so the acceleration \(a\), of such block is: \(a=F/m\), finally the space \(S\) traveled by the block, under the action of force \(F\), namely inside the interval time \((0,4)\), is: \[\Large S = \frac{1}{2}\frac{F}{m}{\tau ^2} = ...?\] where \(\tau=4\) seconds